OpenStudy (anonymous):

K = { R, E, T} and G = { 3, 5, 8, 0, 2}. A code is made up of 2 different symbols from K followed by 4 not necessarily different symbols from G. How many codes are possible?

OpenStudy (anonymous):

Why wouldn't it be 3^2 * 5^4 since you are only picking 2 out of the first 3.

OpenStudy (anonymous):

Sorry, \(3\cdot 2\cdot 5^4=3750\)

OpenStudy (anonymous):

That is what the answer is, but I don't understand why it is 3*2 and not 3^2

OpenStudy (anonymous):

\(3\cdot 2\)

OpenStudy (anonymous):

Thank you!